# Deep Learning based Key Information Extraction from Business Documents: Systematic Literature Review

### v1

@@ 5.2.3. Grid-based methods
As mentioned in section 5.1, this paradigm is less popular in related literature and only 10 approaches make use of such grid structures. The analysis of these is shown in table LABEL:gridtab of the electronic supplementary material. An interesting observation is that most of the identified approaches represent evolutions of existing grid-based systems, in particular based on Chargrid (Katti et al., 2018). The only methods that propose novel concepts are (Katti et al., 2018; Gal et al., 2019; Palm et al., 2019; Wang et al., 2021b). This shows that related research is mostly conducted within a very narrow corridor. Regarding the chosen granularities, a majority of the systems is defined on word level (Denk and Reisswig, 2019; Gal et al., 2019; Palm et al., 2019; Kerroumi et al., 2021; Lin et al., 2021; Do et al., 2023), while three are defined on character level (Katti et al., 2018; Dang and Thanh, 2020; Yeghiazaryan et al., 2022) and one method on token level (Wang et al., 2021b).
@@
Table LABEL:gridtab also lists findings regarding the grid dimensions. Almost all methods define the height and width of the grid based on the pixel counts of the input image. The grids are therefore rectangular, as this corresponds to the rectangular format of corresponding VRDs. (Gal et al., 2019) and (Yeghiazaryan et al., 2022) use a fixed value for the height and width, namely 512 and 336 respectively, therefore resulting in square grids. (Kerroumi et al., 2021) scale down the input dimensions by a factor of 8 in order to reduce the overall model complexity. The work of (Wang et al., 2021b) use a somewhat different approach and employ a dynamic grid structure, in which the height and width dimensions are defined by the distance between the maximum and minimal coordinates of bounding boxes in vertical and horizontal directions, i.e., the span of actual document content. Another distinguishing factor is the feature vector length. Relatively similar values are chosen here, with 256 and 768 being the most common vector sizes. The approach of (Gal et al., 2019) uses the smallest feature set with a vector of length 35. The most complex feature vector is defined by (Palm et al., 2019). In this case, the feature vector has the dimension 4x128x103.
@@
Regarding the design choice, which features are chosen for the individual grid elements, the majority of approaches use text embeddings obtained from different models. A common choice are BERT-based models including BERT, RoBERTa and PhoBERT (Denk and Reisswig, 2019; Lin et al., 2021; Yeghiazaryan et al., 2022; Do et al., 2023). Another common feature type is a 1-hot encoding of the characters (Katti et al., 2018; Palm et al., 2019; Dang and Thanh, 2020), which means that each character in the word corpus is assigned to a specific index value. Only the approaches by (Gal et al., 2019; Lin et al., 2021; Wang et al., 2021b; Do et al., 2023) integrate visual features, which are based on pixel-wise RGB-values, ResNet18, or Swin Transformer embeddings. Another distinguishing feature is the integration of features for background elements of the grid, i.e., all elements that are not overlapped by contents of the document image. In this regard, almost all approaches use an all-zero feature vector. The aim is to clearly differentiate actual content and backgrounds. (Palm et al., 2019) use a sparse tensor and therefore discard background elements entirely. Only the work of (Kerroumi et al., 2021) integrate features for the background elements, namely the RGB-channels of the corresponding coordinates in order to fuse visual information into the architecture.
@@
Most of the methods employ a semantic segmentation (Katti et al., 2018; Denk and Reisswig, 2019; Dang and Thanh, 2020; Kerroumi et al., 2021; Lin et al., 2021; Yeghiazaryan et al., 2022) to perform the KIE task, where each element of the document image is assigned to semantic categories (i.e., fields to be extracted), which results in a segmentation mask of the original document image. Other choices are bounding box regression (Katti et al., 2018; Denk and Reisswig, 2019), where the goal is to predict bounding boxes of semantic entities, which can be helpful to better differentiate individual objects. For example, when processing an invoice, it is helpful to separate each individual line item. This is also the reason why (Katti et al., 2018; Denk and Reisswig, 2019; Yeghiazaryan et al., 2022) combine semantic segmentation with methods for bounding box regression (or line item detection in general). We refer to table LABEL:gridtab for a detailed overview regarding employed KIE methods.
@@ 5.2.4. Other methods
Of the 96 papers analyzed, 6 cannot be assigned to one of the paradigms of KIE methods discussed before. In the following, we briefly present key concepts behind those deviating approaches. The approach by (Majumder et al., 2020) is based on learning representations document snippets and classifying them into fields to be extracted. First, potential candidates are identified for each field based on the data type. The next step is to select the correct candidate for each field. To this end, a representation is obtained for each candidate as well as each field to be extracted. The candidate representation is constructed based on the candidate itself as well as its neighboring segments, utilizing textual and positional information processed by a self-attention mechanism. Finally, a similarity score is calculated for each pair of candidate embedding and field embedding. The similarity scores are then used in a separate module to select the appropriate candidate for each entity, which can be defined in a variety of ways. A trivial method is to select the candidate with the highest similarity score for each entity.
(Wang et al., 2020) use a hierarchical tree-like structure to represent fragments of the document pages, which can be compared to the graph-based methods. Compared to traditional tree architectures, nodes on the same hierarchy level can also be connected with each other. Parent and child nodes of these trees represent key-value pairs. The KIE task is performed by predicting the relations between the individual fragments, i.e., directions of the edges. Obtaining the most likely parent element of each fragment can be ultimately used for identifying fields to be extracted.
@@
(Sarkhel and Nandi, 2021) follow concepts of (Majumder et al., 2020) in the sense that for each field to be extracted, a set of candidate spans is identified first. In order to identify candidate spans for a named entity, multiple detector functions are implemented, to some extent based on domain-knowledge. The authors then employ an adversarial neural network in order to find local contexts of the identified visual spans. These local contexts represent specific fragments of a document image that contain the spans as well as related relevant context (e.g., a larger paragraph). Based on these identified segments, both global and local context vectors are constructed based on textual and visual features. These features are then used by a binary classification model in order to predict, whether a visual span contains a specific field to be extracted or not.
@@
(Shi et al., 2023) propose a novel document representation structure which they refer to as cell-based. This representation is similar to grid-based methods, however the cell-based methodology has no consistent placement of elements in terms of height and width. Instead, the cells are defined depending on the actual document content. For example, different lines or columns in the cell-structure can consist of a different number of elements. The individual cells are also sorted by row and column index respectively, which provides additional information to the KIE system. The obtained cell-based layout is then processed by sequence-based methods such as LayoutLM and therefore follow the typical sequence labeling scenario that respective models use to perform KIE.
@@ 5.3.1. Methodology and setup
Table LABEL:evaltab of the electronic supplementary material shows the findings regarding the presented evaluation setups. In 18 of the analyzed papers, an element-based evaluation method is chosen, which means that elements like the predicted class of a token are compared with the groundtruth class in case of sequence-based approaches. Another common element-based evaluation is to compare the predicted bounding box with the actual bounding box and to identify a match or mismatch based on the overlap of their respective coordinates. 32 of the analyzed approaches incorporate a string-based evaluation setting. In such cases, the extracted textual values are compared with the groundtruth strings to determine, whether a prediction for a given field was correct or not. A string-based evaluation is particularly relevant for assessing the suitability for real-world applications where the extracted texts are used for further processing (e.g., transfer to other information systems). In cases where the authors either did not explicitly specify their chosen method and/or it was not obvious from the manuscript itself, we declared the property as ”n/a”. This was the case for half of the analyzed papers, showing that authors often do not describe their evaluation procedures in great detail.
@@
In almost all of the manuscripts, the authors present the overall performance, i.e., the aggregated results across all fields of interest and all documents of the test set. However, the field-level performance, which shows the performance for each field individually, is only adopted in 35 of the analyzed papers. One aspect that is almost never presented is the performance on the Unknown class. In the context of KIE, Unknowns represent all elements of a document that do not belong to any of the fields to be extracted. This can be helpful in determining the extent to which an approach is able to distinguish irrelevant content from relevant parts of a document image. However, only (Krieger et al., 2021; De Trogoff et al., 2022; Gemelli et al., 2023) explicitly report the evaluation results for this class. In most cases (n=85), the proposed method is compared against existing KIE systems, either by comparing the own results with the evaluation metrics presented in the respective papers or by re-implementation. Thus, it can be stated that a high emphasis is placed on the overall comparability and the legitimization of the own approach. However, especially when authors decide to re-implement existing approaches, comparisons of evaluation results are not always appropriate, since less effort is usually put into training the approaches of third parties.
@@
Considering the employed evaluation metrics, one can see that the F1 score is predominantly used, more specifically in 81 manuscripts. Precision and Recall, on which the F1 score is based, are only explicitly reported in about 25 % of the cases. Interestingly, the usually widely used Accuracy metric is only adopted by (Gal et al., 2019; Palm et al., 2019; Cheng et al., 2020; Gal et al., 2020; Belhadj et al., 2021; Klaiman and Lehne, 2021; Ning et al., 2021; Sarkhel and Nandi, 2021; De Trogoff et al., 2022). One reason for this is that the task of KIE is a multi-class classification problem and that the datasets are usually very unbalanced in terms of the distribution of the fields to be extracted compared to the Unknowns, which strongly distorts the Accuracy metric (Galar et al., 2012). Moreover, it is noticeable that a large majority of the identified evaluation metrics (15 out of 24) are custom-defined and only used in one particular paper, which prevents an adequate benchmark against existing KIE approaches.
@@
Custom datasets are being used in 45 of the analyzed papers and typically represent in-house document collections that are not publicly available. However, in most of the cases when custom datasets are used, the authors also present their results on public benchmark datasets to allow for a comparative evaluation. In addition, there are numerous datasets that were proposed in individual papers, but were not adopted in any other related work. This is the case for 16 of the 26 identified datasets. On average, two datasets are used for evaluation purposes.
@@
Dataset Documents # docs # classes555in terms of entities to extract Language # uses666as part of analyzed approaches Pre-Training?
FUNSD Forms 199 4 eng 43
CORD Receipts 1,000 30 eng 34
SROIE Receipts 973 4 eng 32
IIT-CDIP Lawsuits 6,000,000 / eng 21
EPHOIE Exams 1,494 10 zho 7
RVL-CDIP Lawsuits 400,000 / eng 6
Kleister-NDA NDAs 540 4 eng 5
XFUND Forms 1,393 5 zho,jpn,spa,fra,ita,deu,por 5
DocBank Miscellaneous 500,000 12 eng 4
Table 2 provides an overview of the most common datasets adopted in KIE research. Listed are benchmark datasets as well as pre-training datasets. By far the most common datasets in this regard are FUNSD (Jaume et al., 2019), CORD (Park et al., 2019) and SROIE (Huang et al., 2019), all of which are used by at least a third of the papers. FUNSD includes forms from various domains. CORD and SROIE on the other hand contain photographed receipts, mostly from supermarkets and restaurants. There are discrepancies both in terms of dataset size and in the number of fields/keys to be extracted. The datasets with many hundreds of thousands of documents are typically used for the pre-training of corresponding models and not as an evaluation benchmark. Therefore, these datasets are also widely adopted in other DU tasks. Two of the three most used datasets only aim at extracting four fields of interest, which is a relatively low amount compared to the variety of information corresponding documents usually include. CORD on the other hand includes 30 fields to be extracted, which is also the highest among the listed datasets. The key difference between CORD and SROIE is that in case of former, detailed information including individual line item attributes such as their quantity or unit price need to be extracted, while SROIE only requires to extract aggregated information such as the total price. For the most part, the datasets include English documents. One exception is XFUND (Xu et al., 2022), which specifically intends to investigate multi-lingual capabilities of KIE approaches. Two datasets based on Chinese documents exist, namely EPHOIE (Wang et al., 2021a) and Ticket (Guo et al., 2019), however they are understandably not as widely adopted as most English counterparts that are being used by the international research community.

### v2

@@ 5.2.3. Grid-based methods
As mentioned in section 5.1, this paradigm is less popular in related literature and only 10 approaches make use of such grid structures. This suggest that while it offers an intuitive document representation, grid-based systems struggle with flexibility and generalizability across diverse and complex layouts of VRDs. The analysis of the approaches is shown in table ST6 of the electronic supplementary material. An interesting observation is that most approaches represent evolutions of existing grid-based systems, in particular Chargrid (Katti et al., 2018). This shows that related research is mostly conducted within a very narrow corridor. Regarding the chosen granularities, a majority of the systems is defined on word level, while three are defined on character level and one method even on token level.
@@
Table ST6 also lists findings regarding the grid dimensions. Almost all methods define the height and width of the grid based on the pixel counts of the input image. The grids are therefore rectangular, as this corresponds to the rectangular format of corresponding VRDs. (Gal et al., 2019) and (Yeghiazaryan et al., 2022) use a fixed value for the height and width, namely 512 and 336 respectively, therefore resulting in square grids. (Kerroumi et al., 2021) scale down the input dimensions by a factor of 8 in order to reduce the overall model complexity. The work of (Wang et al., 2021b) uses a somewhat different approach and employs a dynamic grid structure, in which the height and width dimensions are defined by the distance between the maximum and minimal coordinates of bounding boxes in vertical and horizontal directions – i.e., the span of actual document content. Another distinguishing factor is the feature vector length. Relatively similar values are chosen here, with 256 and 768 being the most common vector sizes. The approach of (Gal et al., 2019) uses the smallest feature set with a vector of length 35. The most complex feature vector is defined by (Palm et al., 2019). In this case, the feature vector has the dimension 4×128×103.
@@
Regarding which features are chosen for the individual grid elements, the majority of approaches use text embeddings obtained from different models – mostly BERT-based architectures. Another common feature type is a 1-hot encoding of the characters, which means that each character in the word corpus is assigned to a specific index value. A few approaches also integrate visual features based on pixel-wise RGB-values, ResNet18, or Swin Transformer embeddings. Another distinguishing feature is the integration of features for background elements of the grid, i.e., all elements that are not overlapped by contents of the document image. In this regard, almost all approaches use an all-zero feature vector. The aim is to clearly differentiate actual content and backgrounds. (Palm et al., 2019) use a sparse tensor and therefore discard background elements entirely. Only the work of (Kerroumi et al., 2021) integrate features for the background elements, namely the RGB-channels of the corresponding coordinates in order to fuse visual information into the architecture.
@@
Most of the methods employ semantic segmentation to perform the KIE task, where each element of the document image is assigned to semantic categories (i.e., fields to be extracted), which results in a segmentation mask of the original document image. Other choices are bounding box regression, where the goal is to predict bounding boxes of semantic entities, which can be helpful to better differentiate individual objects. For example, when processing an invoice, it is helpful to separate each individual line item. This is also the reason why a few approaches combine semantic segmentation with methods for bounding box regression (or line item detection in general) (Katti et al., 2018; Denk and Reisswig, 2019; Yeghiazaryan et al., 2022).
5.2.4. LLM-based methods
With the emergence of LLMs across many research areas, their application has also been studied for KIE in the recent past. Multimodal LLMs in particular have accelerated this process, as they can adequately process VRDs based on different modalities. Due to the nature of the models, KIE is typically formulated as a VQA task. That is, the LLMs receive OCR text as input and are prompted with questions like ”What is the value for the {key}?”. All LLM-based approaches are generative by design. A key advantage of these approaches lies in the basis of LLMs: they have been pre-trained on extremely large amounts of data, significantly larger than those used in KIE-specialized sequence-based models (see also section 5.2.1). This allows them to draw on extensive world knowledge, which can result in a higher generalization capability compared to more traditional KIE methods. The analysis of LLM-based approaches is provided in table ST7 of the electronic supplementary material.
More than half of the approaches use an external OCR engine. This shows that robust text recognition remains important and should not be left solely to the LLM itself. This point is underlined by the fact that only (Ye et al., 2023; Feng et al., 2024; Liu et al., 2024) operate purely on the image modality. Almost all methods use open-source LLMs as a backbone, with LLama2 being particularly common. In most cases, a visual encoder is used to integrate the visual information. To this end, three of the approaches use LayoutLMv3 for image encoding. The works vary in terms of image resolution. Some use high resolutions such as 2560×2560, while others use lower resolutions like 224×224 but compensated by extracting up to 20 image crops in order to more precisely capture individual document regions. Here, a trade-off must be made between computational overhead and extraction quality.
Some approaches perform continued pre-training to better adapt the backbone LLMs to VRDs. Tasks used during this phase include text recognition, text spotting, and table understanding. This continued pre-training is often carried out on existing KIE datasets, such as IIT-CDIP. With the exception of (He et al., 2023b), all approaches additionally perform supervised fine-tuning (SFT) for KIE and therefore adjust model weights. The goal is to improve the foundation models’ capabilities of processing VRDs and perform KIE. For this purpose, corresponding instruction-tuning datasets are created, often based on popular benchmark datasets such as CORD. This involves converting the document annotations into question-answer pairs similar to the VQA-style question mentioned before. Besides, some authors also include various other SFT tasks such as image captioning where the LLM is prompted to generate textual captions for input documents. In this regard, very few approaches use parameter-efficient fine-tuning methods such as LoRA. Instead, most train all parameters of the LLMs during SFT. This indicates that in order to properly fine-tune the foundation models to VRDs and KIE, it is required to update the weights of the majority of the architecture. However, there is no work that analyzes this aspect in depth.
The approaches also differ in how the different modalities are fed to the LLMs. For example, (He et al., 2023b) pass OCR data to the LLM and prompt it to obtain the corresponding class. An example format is: ”Q3:{text: ’TO:’, Box:[102 345 129 359]}…, What are the labels for these texts?”. Incorporating layout information explicitly by adding a list representation of the bounding boxes into the prompt like above is a straightforward and commonly used approach. However, it is associated with limitations. Using this method, a large part of the input sequence is therefore occupied by these tokens, which reduces the available context length for the actual VRD content. Also, the backbone LLMs are typically not (pre-)trained with such prompt formats. This can lead to the model struggling to properly reason about the layout information.
(Minouei et al., 2024) take a different approach by converting documents into an HTML structure and passing it to the LLM in order to obtain a structured JSON output with the extracted information. The authors argue that LLMs are well-suited to processing HTML, as a large portion of their pre-training data consists of HTML. Besides, the common procedure of LLM-based approaches is to obtain image encodings and pass them to the LLM together with the textual prompt to get the extracted value for the respective field. Some works include all fields to be extracted in one prompt (e.g., (Minouei et al., 2024)), while others only consider one field per call (e.g., (Ye et al., 2023)). The latter requires multiple calls per document, which can increase runtime, but may allow the LLM to better focus on one field and improve extraction quality per field. Authors also experiment with different prompt templates. For example, KIE is sometimes framed as a multiple-choice task: {document} What is ‘{value}’ in the document? Possible choices: {keys}, where {keys} is a randomly selected subset of key names from the dataset (Yoon et al., 2024). The idea is to leverage the LLM’s prior knowledge of field names to better identify the correct value. In general, most works follow similar prompt formulations. However, (Zmigrod et al., 2024) shows that it is advantageous to create instruction datasets with heterogeneous prompts, as this increases robustness for different layouts of VRDs.
@@ 5.2.5. Other methods
Of the 130 papers analyzed, seven cannot be assigned to any of the paradigms of KIE methods discussed before. In the following, we briefly present key concepts behind those deviating approaches. The approach by (Majumder et al., 2020) is based on learning representations of document snippets and classifying them into fields to be extracted. First, potential candidates are identified for each field based on the data type. The next step is to select the correct candidate for each field. To this end, a representation is obtained for each candidate as well as each field to be extracted. The candidate representation is constructed based on the candidate itself as well as its neighboring segments, utilizing textual and positional information processed by a self-attention mechanism. Finally, a similarity score is calculated for each pair of candidate embedding and field embedding. The similarity scores are then used in a separate module to select the appropriate candidate for each entity, which can be defined in a variety of ways. A trivial method is to select the candidate with the highest similarity score for each entity.
(Wang et al., 2020) use a hierarchical tree-like structure to represent fragments of the document pages, similar to the graph-based methods. Compared to traditional tree architectures, nodes on the same hierarchy level can also be connected with each other. Parent and child nodes of these trees represent key-value pairs. The KIE task is performed by predicting the relations between the individual fragments, i.e., directions of the edges. Obtaining the most likely parent element of each fragment can be ultimately used for identifying fields to be extracted.
@@
(Sarkhel and Nandi, 2021) follow concepts of (Majumder et al., 2020) in the sense that for each field to be extracted, a set of candidate spans is identified first. In order to identify candidate spans for a named entity, multiple detector functions are implemented, to some extent based on domain knowledge. The authors then employ an adversarial neural network in order to find local contexts of the identified visual spans. These local contexts represent specific fragments of a document image that contain the spans as well as related relevant context (e.g., a larger paragraph). Based on these identified segments, both global and local context vectors are constructed with textual and visual features. These features are then used by a binary classification model in order to predict, whether a visual span contains a specific field to be extracted or not.
@@
(Tang et al., 2023) propose a novel document representation structure which they refer to as cell-based. This representation is similar to grid-based methods, however the cell-based methodology has no consistent placement of elements in terms of height and width. Instead, the cells are defined depending on the actual document content. For example, different lines or columns in the cell-structure can consist of a different number of elements. The individual cells are also sorted by row and column index respectively, which provides additional information to the KIE system. The obtained cell-based layout is then processed by sequence-based methods such as LayoutLM and therefore follow the typical sequence labeling scenario that respective models use to perform KIE.
Instead of processing entities sequentially or constructing elaborate graph structures, the approach by (Lin et al., 2024) constructs a token pair representations matrix based on multi-modal features from a pre-trained encoder. It then jointly generates three relation matrices: line extraction (to determine the boundaries of text lines), line grouping (to merge related lines that form multi-line entities), and entity linking (to connect keys with their corresponding values). These predictions are then combined to ultimately obtain extracted key-value pairs.
@@ 5.3.1. Methodology and setup
Table ST8 of the electronic supplementary material shows the findings regarding the presented evaluation setups. A consistent picture throughout is that authors often do not describe their evaluation procedures in great detail and it is also not always obvious from the manuscript itself. This is the case for around half of the analyzed papers. We declared affected properties as ”n/a”. In 19 of the analyzed papers, an element-based evaluation method is chosen, which means that elements like the predicted class of a token are compared with the groundtruth class in case of sequence-based approaches. Another common element-based evaluation is to compare the predicted bounding box with the actual bounding box and to identify a match or mismatch based on the overlap of their respective coordinates. 45 of the analyzed approaches incorporate a string-based evaluation setting. In such cases, the extracted textual values are compared with the groundtruth strings to determine, whether a prediction for a given field was correct or not. A string-based evaluation is particularly relevant for assessing the suitability for real-world applications where the extracted texts are used for further processing (e.g., transfer to other information systems).
@@
In almost all of the manuscripts, the authors present the overall performance, i.e., the aggregated results across all fields of interest and all documents of the test set. However, the field-level performance, which shows the performance for each field individually, is only adopted in around 30% of the analyzed papers. One aspect that is almost never presented, is the performance on the Unknown class. In the context of KIE, Unknowns represent all elements of a document that do not belong to any of the fields to be extracted. This can be helpful in determining the extent to which an approach is able to distinguish irrelevant content from relevant parts of a document image. In most cases, the proposed method is compared against existing KIE systems, either by comparing the own results with the evaluation metrics presented in the respective papers or by re-implementation. Thus, it can be stated that a high emphasis is placed on the overall comparability and the legitimization of the own approach. However, especially when authors decide to re-implement existing approaches, comparisons of evaluation results are not always appropriate, since less effort is usually put into training the approaches of third parties.
@@
Considering the employed evaluation metrics, one can see that the F1 score is used most of the time. Precision and Recall, with the F1 score being the harmonic mean of these two metrics, are only explicitly reported in about 25% of the cases. Interestingly, the usually widely used Accuracy metric is only presented by a few authors. One reason for this is that KIE datasets are usually very unbalanced in terms of label distribution, which strongly distorts the overall Accuracy metric (Galar et al., 2012) – making it less representative for these scenarios. Nonetheless, given that KIE is usually seen as a multi-class classification problem, it is not surprising that authors adopt metrics such as Precision, Recall and F1, which are very common for such tasks. As mentioned before, many authors present a string-based evaluation. To this end, various different metrics are used, e.g., Word Accuracy Rate, which counts the number of substitutions, insertions and deletions between groundtruth and predictions (Kerroumi et al., 2021). Moreover, it is noticeable that a large majority of the identified evaluation metrics are custom-defined and only used in one particular paper, which prevents an adequate benchmark against existing KIE approaches.
@@
Custom datasets are being used in 54 of the analyzed papers and typically represent in-house document collections that are not publicly available. However, in around 80% of the cases when custom datasets are used, the authors additionally present their results on public benchmark datasets to allow for a comparison with other KIE systems. In addition, there are numerous datasets that were proposed in individual papers, but were not adopted in any other subsequent work. This indicates again that KIE literature strongly focuses on a small set of benchmark datasets.
@@
Dataset Documents # docs # classes444in terms of entities to extract Language # uses555as part of analyzed approaches Pre-Training?
FUNSD Forms 199 4 eng 61
CORD Receipts 1,000 30 eng 52
SROIE Receipts 973 4 eng 45
IIT-CDIP Lawsuits 6,000,000 / eng 24
XFUND Forms 1,393 4 zho,jpn,spa,fra,ita,deu,por 15
EPHOIE Exams 1,494 10 zho 9
RVL-CDIP Lawsuits 400,000 / eng 7
Kleister-NDA NDAs 540 4 eng 6
Kleister-Charity Reports 2,778 8 eng 6
Table 2 provides an overview of the most common datasets adopted in KIE research. Listed are benchmark datasets as well as pre-training datasets. By far the most common datasets in this regard are FUNSD (Jaume et al., 2019), CORD (Park et al., 2019) and SROIE (Huang et al., 2019), all of which are used by at least a third of the papers. FUNSD includes forms from various domains. CORD and SROIE on the other hand contain photographed receipts, mostly from supermarkets and restaurants. There are discrepancies both in terms of dataset size and in the number of keys to be extracted. The datasets with many hundreds of thousands of documents are typically used for the pre-training of corresponding models and not as an evaluation benchmark. Therefore, these datasets are also widely adopted in other DU tasks. Two of the three most used datasets only aim at extracting four fields of interest, which is a relatively low amount compared to the variety of information corresponding documents usually include. CORD on the other hand includes 30 fields to be extracted, which is also the highest among the listed datasets. The key difference between CORD and SROIE is that in case of former, detailed information including individual line item attributes such as their quantity or unit price need to be extracted, while SROIE only requires to extract aggregated information such as the total price. For the most part, the datasets include English documents. One exception is XFUND (Xu et al., 2022), which specifically intends to investigate multi-lingual capabilities of KIE approaches. Some popular Chinese datasets exist, namely EPHOIE and Ticket, however they are understandably not as widely adopted as most English counterparts that are being used by the international research community.